OpenStudy (anonymous):

A chemist mixes some 70% solution with some 40% solution to obtain 120 gallons of 50% solution. Find the number of gallons of each solution. I've stated the variables which is x= the number of gallons for 70% and y= the number of gallons for 40%.

OpenStudy (aum):

"x" gallons of 70% solution is mixed with "y" gallons of 40% solution to form 120 gallons of 50% solution. x + y = 120 ----- (1) 0.7x + 0.4y = 120*0.5 = 60 ----- (2) Solve for x and y.